### v1

# Full asymptotic expansions of the Humbert function $\Phi_{1}$

Abstract
We derive full asymptotic expansions for the Humbert function $\Phi_{1}$ in different limiting regimes of its variables. Our derivation employs various asymptotic methods and relies on key transformation formulae established by Erdélyi (1940), and Tuan and Kalla (1987). The efficiency of our asymptotic results are also illustrated through two applications: (1) analytic continuations of Saran’s function $F_{M}$, and (2) two limits arising in the study of the $1D$ Glauber-Ising model. Finally, some promising directions for future research are highlighted.
2020 Mathematics Subject Classification: 33C65, 41A60, 82C20.
Keywords: asymptotic expansions, Humbert function, Glauber-Ising model.
1.  Introduction
The Humbert function $\Phi_{1}$ was first introduced by Humbert [14] and is defined as
$\Phi_{1}[a,b;c;x,y]=\sum_{m,n=0}^{\infty}\frac{\left(a\right)_{m+n}\left(b\right)_{m}}{\left(c\right)_{m+n}}\frac{x^{m}}{m!}\frac{y^{n}}{n!},\quad|x|<1,|y|<\infty,$
where $a,b\in\mathbb{C},\,c\in\mathbb{C}\setminus\mathbb{Z}_{\leqslant 0}$ and $\left(\alpha\right)_{n}=\frac{\Gamma(\alpha+n)}{\Gamma(\alpha)}$. This function has the series representation [14, p. 75]
$\Phi_{1}[a,b;c;x,y]=\sum_{n=0}^{\infty}\frac{\left(a\right)_{n}}{\left(c\right)_{n}}\,_{2}F_{1}\left[\begin{matrix}a+n,b\\ c+n\end{matrix};x\right]\frac{y^{n}}{n!}.$ (1.1)
Applying the Euler transformation [18, Eq. (15.8.1)]
${}_{2}F_{1}\left[\begin{matrix}a,b\\ c\end{matrix};x\right]=\left(1-x\right)^{-b}\,_{2}F_{1}\left[\begin{matrix}c-a,b\\ c\end{matrix};\frac{x}{x-1}\right],$ (1.2)
one may obtain the trivial estimate
${}_{2}F_{1}\left[\begin{matrix}a+s,b\\ c+s\end{matrix};x\right]=\mathcal{O}(1),\quad|c+s|\to\infty,$ (1.3)
which then extends $\Phi_{1}$ to the region
$\mathbb{D}=\left\{(x,y)\in\mathbb{C}^{2}:x\neq 1,\ \left|\arg(1-x)\right|<\pi,\ |y|<\infty\right\}.$ (1.4)
Over the past century, the Humbert function $\Phi_{1}$ has been extensively studied in both mathematics and physics. In his seminal 1939-1940 works [7, 8], Erdélyi derived fundamental identities for the bivariate confluent hypergeometric functions $\Phi_{1},\Phi_{2},\Psi_{1}$ and $\Gamma_{1}$ by systematically investigating their defining systems of partial differential equations. It should be noted that, building upon Erdélyi’s foundational work, Shimomura [21, 22] derived complete asymptotic expansions of the Humbert function $\Phi_{2}$ in cases when one or two variables become large. In 2014, Del Punta et al. [5] showed that solutions for the pure Coulomb potential with a driven term involving either Slater-type or Laguerre-type orbitals can be expressed as linear combinations of $\Phi_{1}$ functions. They subsequently obtained the asymptotic expansions of those solutions through formal analysis of the large-$y$ behavior of $\Phi_{1}$. In 2022, Brychkov et al. [4] expressed the cumulative distribution function of the non-central $\chi_{\nu}^{\prime 2}(\lambda)$ distribution in closed form via specific, equal-parameter $\Phi_{1}$ function. They also derived connection formulae between specific Humbert functions $\Phi_{1}$ and $\Phi_{3}$.
To the best of our knowledge, prior to this paper, there has been no investigation in the existing literature regarding the full asymptotic expansions of $\Phi_{1}$. In this paper, we systematically study the asymptotic behaviors of $\Phi_{1}$ under different limiting cases. Table 1 summarizes our main results and the tools employed.
Function Limiting Case Tools Results
$\Phi_{1}[x,y]$ $x\to\infty$ Series manipulation technique Theorem 2.1
$\Phi_{1}[x,y]$ $x\to 1$ Connection formula Theorem 2.2
$\Phi_{1}[x,y]$ $\Re(y)\to-\infty$ Mellin-Barnes integral Theorem 2.3
$\Phi_{1}[x,y]$ $\Re(y)\to+\infty$ Kummer transformation Corollary 2.5
$\Phi_{1}[x,y]$ $x\to\infty,y\to\infty$ Connection formula Theorems 2.7 and 2.8
$\Phi_{1}\bigl{[}x,\frac{\eta}{x}\bigr{]}$ $x\to\infty$ Series manipulation technique Theorem 2.9
$\Phi_{1}\bigl{[}\frac{\eta}{y},y\bigr{]}$ $y\to\infty$ Uniformity approach Theorem 2.10
Table 1: Main results and the tools employed.
The remaining part of the paper is structured as follows. In Section 2, we derive the asymptotic expansions of $\Phi_{1}$ by applying various asymptotic methods and using novel transformation formulae established by Erdélyi [8], and Tuan and Kalla [25], which prove particularly crucial for investigating Horn’s functions (see [9, pp. 224-227]). Section 3 provides two practical examples, including analytic continuations of Saran’s triple hypergeometric function $F_{M}$ and two limits appearing in the study of $1D$ Glauber-Ising model, to demonstrate the applicability of our asymptotic results. The paper concludes in Section 4 with a brief discussion on future study.
Notation. By $f(n,z)=\mathcal{O}(a_{n}g(z))\,(z\in\Omega)$, we mean that there exists a constant $K>0$ independent of $n$ and $z$ such that
$|f(n,z)|\leqslant K|a_{n}||g(z)|,\quad n\in\mathbb{Z}_{\geqslant 0},\,z\in\Omega.$
Moreover, the generalized hypergeometric function ${}_{p}F_{q}$ is defined by
${}_{p}F_{q}\left[\begin{matrix}a_{1},\cdots,a_{p}\\ b_{1},\cdots,b_{q}\end{matrix};z\right]\equiv{}_{p}F_{q}[a_{1},\cdots,a_{p};b_{1},\cdots,b_{q};z]:=\sum_{n=0}^{\infty}\frac{\left(a_{1}\right)_{n}\cdots\left(a_{p}\right)_{n}}{\left(b_{1}\right)_{n}\cdots\left(b_{q}\right)_{n}}\frac{z^{n}}{n!},$ (1.5)
where $a_{1},\cdots,a_{p}\in\mathbb{C}$ and $b_{1},\cdots,b_{q}\in\mathbb{C}\setminus\mathbb{Z}_{\leqslant 0}$.
2.  Asymptotics of $\Phi_{1}$
In this section, we deduce the asymptotic behaviors of $\Phi_{1}[x,y]$ for $x$ and $y$ in five different regimes: (i) $x\to\infty$; (ii) $x\to 1$; (iii) $y\to\infty$; (iv) $x\to\infty,\,y\to\infty$; and (v) $x$ or $y$ small, $xy$ fixed.
Remarkably, the singular behavior of $\Phi_{1}$ in Regime (v) differs significantly from the asymptotic behaviors exhibited in Regimes (i)-(iv). We refer to [12] for a comprehensive treatment of this singular case.
2.1.  Regime (i): $x\to\infty$
The following series representation illustrates the full asymptotics of $\Phi_{1}[x,y]$ for large $x$.
Theorem 2.1.
Assume that $a,b\in\mathbb{C},\,c\in\mathbb{C}\setminus\mathbb{Z}_{\leqslant 0}$ and $a-b\in\mathbb{C}\setminus\mathbb{Z}$. Then
$\begin{split}\Phi_{1}[a,b;c;x,y]={}&\frac{\Gamma(c)\Gamma(b-a)}{\Gamma(b)\Gamma(c-a)}\left(-x\right)^{-a}\sum_{k=0}^{\infty}{}_{1}F_{1}\left[\begin{matrix}-k\\ c-a-k\end{matrix};-y\right]\frac{\left(a\right)_{k}\left(a-c+1\right)_{k}}{\left(a-b+1\right)_{k}}\frac{x^{-k}}{k!}\\ &+\frac{\Gamma(c)\Gamma(a-b)}{\Gamma(a)\Gamma(c-b)}\left(-x\right)^{-b}\sum_{k=0}^{\infty}{}_{1}F_{1}\left[\begin{matrix}c-b-k\\ a-b-k\end{matrix};y\right]\frac{\left(b\right)_{k}\left(b-a+1\right)_{k}}{\left(b-c+1\right)_{k}}\frac{x^{-k}}{k!}\end{split}$ (2.1)
holds for $\left|\arg(-x)\right|<\pi$, $|x|>1$ and $|y|<\infty$.
Proof..
Using the connection formula for ${}_{2}F_{1}$ ([18, Eq. (15.8.2)])
${}_{2}F_{1}\left[\begin{matrix}a,b\\ c\end{matrix};z\right]={}$ $\frac{\Gamma(c)\Gamma(b-a)}{\Gamma(b)\Gamma(c-a)}\left(-z\right)^{-a}\,_{2}F_{1}\left[\begin{matrix}1-c+a,a\\ 1-b+a\end{matrix};\frac{1}{z}\right]$
$+\frac{\Gamma(c)\Gamma(a-b)}{\Gamma(a)\Gamma(c-b)}\left(-z\right)^{-b}\,_{2}F_{1}\left[\begin{matrix}1-c+b,b\\ 1-a+b\end{matrix};\frac{1}{z}\right],\ \ |\arg(-z)|<\pi.$
one may get from (1.1) that
$\Phi_{1}[a,b;c;x,y]=\frac{\Gamma(c)\Gamma(b-a)}{\Gamma(b)\Gamma(c-a)}\left(-x\right)^{-a}U_{1}(x,y)+\frac{\Gamma(c)\Gamma(a-b)}{\Gamma(a)\Gamma(c-b)}\left(-x\right)^{-b}U_{2}(x,y),$ (2.2)
where
$U_{1}(x,y)$ $=\sum_{n=0}^{\infty}\frac{\left(a\right)_{n}}{\left(a-b+1\right)_{n}n!}\,_{2}F_{1}\left[\begin{matrix}a+n,a-c+1\\ a-b+1+n\end{matrix};\frac{1}{x}\right]\left(\frac{y}{x}\right)^{n},$
$U_{2}(x,y)$ $=\sum_{n=0}^{\infty}\frac{\left(a-b\right)_{n}}{\left(c-b\right)_{n}}\,_{2}F_{1}\left[\begin{matrix}b,b-c+1-n\\ b-a+1-n\end{matrix};\frac{1}{x}\right]\frac{y^{n}}{n!}.$
The series $U_{1}$ is a special case of $\Phi_{1}$. In fact, we obtain from (1.1) that
$U_{1}(x,y)$ $=\Phi_{1}\left[a,a-c+1;a-b+1;\frac{1}{x},\frac{y}{x}\right]$
$=\sum_{m,n=0}^{\infty}\frac{\left(a\right)_{m+n}\left(a-c+1\right)_{m}}{\left(a-b+1\right)_{m+n}m!\,n!}\frac{y^{n}}{x^{m+n}},\quad|x|>1,|y|<\infty.$ (2.3)
Further calculation gives
$U_{1}(x,y)$ $=\sum_{k=0}^{\infty}\frac{\left(a\right)_{k}}{\left(a-b+1\right)_{k}}\frac{x^{-k}}{k!}\sum_{n=0}^{k}\left(-k\right)_{n}\left(a-c+1\right)_{k-n}\frac{y^{n}}{m!}$
$=\sum_{k=0}^{\infty}{}_{1}F_{1}\left[\begin{matrix}-k\\ c-a-k\end{matrix};-y\right]\frac{\left(a\right)_{k}\left(a-c+1\right)_{k}}{\left(a-b+1\right)_{k}}\frac{x^{-k}}{k!}.$
The series $U_{2}$ can be expressed in the notation of double Srivastava-Daoust function [23], namely,
$U_{2}(x,y)$ $=\sum_{m,n=0}^{\infty}\frac{\left(b\right)_{m}\left(b-a+1\right)_{m-n}}{\left(b-c+1\right)_{m-n}m!\,n!}\frac{y^{n}}{x^{m}},\quad|x|>1,|y|<\infty$ (2.4)
[…]
In addition, we have
$U_{2}(x,y)$ $=\sum_{m,n=0}^{\infty}\frac{\left(b\right)_{m}\left(b-a+1\right)_{m}}{\left(b-c+1\right)_{m}}\frac{x^{-m}}{m!}\cdot\frac{\left(c-b-m\right)_{n}}{\left(a-b-m\right)_{n}}\frac{y^{n}}{n!}$
$=\sum_{m=0}^{\infty}{}_{1}F_{1}\left[\begin{matrix}c-b-m\\ a-b-m\end{matrix};y\right]\frac{\left(b\right)_{m}\left(b-a+1\right)_{m}}{\left(b-c+1\right)_{m}}\frac{x^{-m}}{m!}.$
A combination of the above formulas gives (2.1). ∎
2.2.  Regime (ii): $x\to 1$
The singularity of $\Phi_{1}[x,y]$ at $x=1$ is determined by the connection formula below.
Theorem 2.2 ([25, Eq. (32)]).
Assume that $a,b\in\mathbb{C},\,c\in\mathbb{C}\setminus\mathbb{Z}_{\leqslant 0}$ and $a+b-c\notin\mathbb{Z}$. Then for $(x,y)\in\mathbb{D}$,
$\Phi_{1}[a,b;c;x,y]={}$ $\frac{\Gamma(c)\Gamma(c-a-b)}{\Gamma(c-a)\Gamma(c-b)}\Psi_{1}[a,b;a+b-c+1,c-b;1-x,y]$
$+\frac{\Gamma(c)\Gamma(a+b-c)}{\Gamma(a)\Gamma(b)}\left(1-x\right)^{c-a-b}\Psi_{1}[c-b,c-a;c-a-b+1,c-b;1-x,y].$
2.3.  Regime (iii): $y\to\infty$
We first establish the asymptotic expansion of $\Phi_{1}[x,y]$ as $y\to\infty$ in the left half-plane by appealing to its Mellin-Barnes integral representation.
Theorem 2.3.
(i) Let $a,b\in\mathbb{C},\,c\in\mathbb{C}\setminus\mathbb{Z}_{\leqslant 0},\,\delta\in\bigl{(}0,\frac{\pi}{2}\bigr{]}$ and
$\mathbb{V}_{\Phi_{1}}=\left\{(x,y)\in\mathbb{C}^{2}:x\neq 1,\,y\neq 0,\,\left|\arg(1-x)\right|<\pi,\,\left|\arg(-y)\right|\leqslant\frac{\pi}{2}-\delta\right\}.$
Then
$\Phi_{1}[a,b;c;x,y]$ $=\frac{1}{2\pi\mathrm{i}}\frac{\Gamma(c)}{\Gamma(a)}\int_{\mathfrak{L}_{\mathrm{i}\sigma\infty}}{}_{2}F_{1}\left[\begin{matrix}a+s,b\\ c+s\end{matrix};x\right]\frac{\Gamma(a+s)}{\Gamma(c+s)}\Gamma(-s)\left(-y\right)^{s}\mathrm{d}s,$ (2.5a)
where the path $\mathfrak{L}_{\mathrm{i}\sigma\infty}$, starting at $\sigma-\mathrm{i}\infty$ and ending at $\sigma+\mathrm{i}\infty$, is a vertical line intended if necessary to separate the poles of $\Gamma(a+s)$ from the poles of $\Gamma(-s)$.
(ii) Let $a,b\in\mathbb{C}$ and $c,c-a\in\mathbb{C}\setminus\mathbb{Z}_{\leqslant 0}$. When $(x,y)\in\mathbb{V}_{\Phi_{1}}$ and $y\to\infty$, we have
$\Phi_{1}[a;c,c^{\prime};x,y]$ $\sim\frac{\Gamma(c)}{\Gamma(c-a)}\sum_{n=0}^{\infty}{}_{2}F_{1}\left[\begin{matrix}-n,b\\ c-a-n\end{matrix};x\right]\frac{\left(a\right)_{n}(a-c+1)_{n}}{n!}\left(-y\right)^{-a-n}.$ (2.5b)
Proof..
(i) The proof is divided into two steps. Firstly, we show the integral (2.5a) exists for $(x,y)\in\mathbb{V}_{\Phi_{1}}$. Secondly, we prove by applying Cauchy’s residue theorem that the integral is equal to $\Phi_{1}$ when $x$ and $y$ are appropriately restricted. Then by analytic continuation, (2.5a) is valid throughout $\mathbb{V}_{\Phi_{1}}$.
Step 1. The integral (2.5a) formally follows from (1.1) by means of Ramanujan’s master theorem [2, p. 107] and the inverse Mellin transform. To check the validity, write
$\Phi(s):={}_{2}F_{1}\left[\begin{matrix}a+s,b\\ c+s\end{matrix};x\right]\frac{\Gamma(a+s)}{\Gamma(c+s)}\Gamma(-s)\left(-y\right)^{s}.$ (2.6)
Recall the estimate [11, Eq. (3.4)]: as $t\to\pm\infty$,
$\left|\frac{\Gamma(a+\sigma+\mathrm{i}t)}{\Gamma(c+\sigma+\mathrm{i}t)}\Gamma(-\sigma-\mathrm{i}t)\left(-y\right)^{\sigma+\mathrm{i}t}\right|=\mathcal{O}\left(\left|t\right|^{\Re(a-c)-\sigma-\frac{1}{2}}\mathrm{e}^{-\left(\frac{\pi}{2}\pm\arg(-y)\right)|t|}\right).$
Applying (1.3), we obtain that as $t\to\pm\infty$,
$\Phi(s)=\mathcal{O}\left(\left|t\right|^{\Re(a-c)-\sigma-\frac{1}{2}}\mathrm{e}^{-\left(\frac{\pi}{2}\pm\arg(-y)\right)|t|}\right).$ (2.7)
Hence when $\left|\arg(-y)\right|\leqslant\frac{\pi}{2}-\delta$, the integrand $\Phi(s)$ decays exponentially as $|s|\to\infty$ along the path $\mathfrak{L}_{\mathrm{i}\sigma\infty}$. Therefore, the integral in (2.5a) is certainly convergent in $\mathbb{V}_{\Phi_{1}}$.
Step 2. Now show that in a region $\widetilde{\mathbb{V}}_{\Phi_{1}}\subset\mathbb{V}_{\Phi_{1}}$, the integral in (2.5a) implies the series representation (1.1). Define
[…]
We close the path of integration by the semi-circle $\mathfrak{C}$, that is parameterized by $s=(N+1/2)\mathrm{e}^{\mathrm{i}\theta}$ $(|\theta|\leqslant\pi/2)$, and then we let $N\to\infty$ through integer values. Recall the estimate [11, p. 7]
$\left|\frac{\Gamma(a+s)}{\Gamma(c+s)}\Gamma(-s)\left(-y\right)^{s}\right|=\left(N^{\Re(a-c)-\frac{1}{2}-\left(N+\frac{1}{2}\right)\cos\theta}\mathrm{e}^{\left(N+\frac{1}{2}\right)(\delta_{1}\cos\theta+\delta_{2}|\sin\theta|)}\right),\quad s\to\infty,\,s\in\mathfrak{C},$
where $\delta_{1}:=1+\log|y|<-1$ and $\delta_{2}:=|\theta|+\left|\arg(-y)\right|-\pi\leqslant-\delta<0$. In view of this and (1.3), the integral on the semi-circle $\mathfrak{C}$ tends to zero as $N\to\infty$. Finally, we get the series representation (1.1) from Cauchy’s residue theorem and the assertion that ([3, p. 7])
$\underset{s=n}{\operatorname*{Res}}\ \Gamma(-s)=\frac{\left(-1\right)^{n-1}}{n!}.$
(ii) Let $I$ denote the integral in the right-hand side of (2.5a), and let $\Phi(s)$ denote the integrand given by (2.6). Choose the positive integer $M\geqslant\max\{1,\Re(-a)\}$ and shift the integration contour to the left (which is permissible on account of the exponential decay of the integrand).
Note that $s=-a-n\ (n\in\mathbb{Z}_{\geqslant 0})$ are the only poles of $\Phi(s)$. Therefore,
$\underset{s=-a-n}{\operatorname*{Res}}\Phi(s)={}_{2}F_{1}\left[\begin{matrix}-n,b\\ c-a-n\end{matrix};x\right]\frac{\Gamma(a+n)}{\Gamma(c-a-n)}\frac{(-1)^{n}}{n!}\left(-y\right)^{-a-n}.$
By Cauchy’s residue theorem, we get
$I$ $=\frac{\Gamma(c)}{\Gamma(a)}\sum_{n=0}^{M}{}_{2}F_{1}\left[\begin{matrix}-n,b\\ c-a-n\end{matrix};x\right]\frac{\Gamma(a+n)}{\Gamma(c-a-n)}\frac{(-1)^{n}}{n!}\left(-y\right)^{-a-n}+\underbrace{\frac{1}{2\pi\mathrm{i}}\frac{\Gamma(c)}{\Gamma(a)}\int_{C_{M}}\Phi(s)\mathrm{d}s}_{\text{denoted by }R_{M}(y)}$
$=\frac{\Gamma(c)}{\Gamma(c-a)}\sum_{n=0}^{M}{}_{2}F_{1}\left[\begin{matrix}-n,b\\ c-a-n\end{matrix};x\right]\frac{(a)_{n}(a-c+1)_{n}}{n!}\left(-y\right)^{-a-n}+R_{M}(y),$
where $C_{M}$ denotes the vertical line $\Re(s)=\Re(-a)-M-1/2$. On account of (2.7), we find that the integral $R_{M}(y)$ is convergent and that $R_{M}(y)=\mathcal{O}_{M}(|y|^{-\Re(a)-M-\frac{1}{2}})$, which completes the proof. ∎
Remark 2.4.
If $b\in\mathbb{C},\,c\in\mathbb{C}\setminus\mathbb{Z}_{\leqslant 0}$ and $a=c+m\,(m\in\mathbb{Z}_{\geqslant 0})$, use of the Kummer transformation [8, p. 348]
$\Phi_{1}[a,b;c;x,y]=\mathrm{e}^{y}\left(1-x\right)^{-b}\Phi_{1}\left[c-a,b;c;\frac{x}{x-1},-y\right]$ (2.8)
and (1.1) yields the reduction formula
$\Phi_{1}[c+m,b;c;x,y]=\mathrm{e}^{y}\left(1-x\right)^{-b}\sum_{n=0}^{m}{}_{2}F_{1}\left[\begin{matrix}n-m,b\\ c+n\end{matrix};\frac{x}{x-1}\right]\binom{m}{n}\frac{y^{n}}{\left(c\right)_{n}}.$
Hence, it follows from (1.2) that
$\Phi_{1}[c+m,b;c;x,y]=\mathrm{e}^{y}\sum_{n=0}^{m}{}_{2}F_{1}\left[\begin{matrix}c+m,b\\ c+n\end{matrix};x\right]\binom{m}{n}\frac{y^{n}}{\left(c\right)_{n}}.$
Asymptotics of $\Phi_{1}[x,y]$ as $y\to\infty$ in the right half-plane then follows from (2.5b) and (2.8).
Corollary 2.5.
Assume that $b\in\mathbb{C}$ and $a,c\in\mathbb{C}\setminus\mathbb{Z}_{\leqslant 0}$. When $(x,-y)\in\mathbb{V}_{\Phi_{1}}$ and $y\to\infty$, we have
$\Phi_{1}[a,b;c;x,y]\sim\frac{\Gamma(c)}{\Gamma(a)}\mathrm{e}^{y}\sum_{n=0}^{\infty}{}_{2}F_{1}\left[\begin{matrix}a,b\\ a-n\end{matrix};x\right]\frac{\left(1-a\right)_{n}(c-a)_{n}}{n!}y^{a-c-n}.$
Remark 2.6.
If $a=-m\,(m\in\mathbb{Z}_{\leqslant 0}),\,b\in\mathbb{C}$ and $c\in\mathbb{C}\setminus\mathbb{Z}_{\leqslant 0}$, the series in (1.1) and thus $\Phi_{1}[-m,b;c;x,y]$ reduce to a finite sum involving Gauss hypergeometric function ${}_{2}F_{1}[x]$.
2.4.  Regime (iv): $x\to\infty,\,y\to\infty$
The full asymptotics of $\Phi_{1}[x,y]$ for large values of $x$ and $y$ follows from the identity [8, p. 348]
$\Phi_{1}[a,b;c;x,y]=\left(1-x\right)^{c-a-b}\mathrm{e}^{\frac{y}{x}}\,\Psi_{1}\left[c-b,c-a;c,c-b;x,\frac{x-1}{x}y\right]$
and from asymptotics of the Humbert function $\Psi_{1}$ for two large arguments (see [13, Theorems 1.2 and 1.3]). We formulate the precise asymptotic expansions in the following two theorems.
Theorem 2.7.
Assume that $c,c-b\in\mathbb{C}\setminus\mathbb{Z}_{\leqslant 0}$ and $b,a-b\in\mathbb{C}\setminus\mathbb{Z}$. Set $\beta=-\frac{y}{x}$. Then, under the condition
$x\to\infty,\,\frac{x-1}{x}y\to+\infty,\quad\left|\arg(1-x)\right|<\pi,\quad 0<\beta_{1}\leqslant|\beta|\leqslant\beta_{2}<\infty,$
the function $\Phi_{1}\equiv\Phi_{1}[a,b;c;x,y]$ admits the asymptotic expansion
$\Phi_{1}\sim\frac{\Gamma(c)}{\Gamma(a)}\beta^{a-c}\left(1-x\right)^{a-b-c}\mathrm{e}^{y}\sum_{k=0}^{\infty}a_{k}^{*}\left(x,y\right)y^{-k},$
where $a_{0}^{*}(x,y)=1$ and in general, for any $k\in\mathbb{Z}_{\geqslant 0}$,
$a_{k}^{*}(x,y)=\sum_{j=0}^{k}\frac{\left(c-a\right)_{j}\left(c-a\right)_{k-j}\left(j+b-a+1\right)_{k-j}}{j!\left(k-j\right)!}\,_{3}F_{2}\left[\begin{matrix}-j,j-k,c-1\\ c-a,a-b-k\end{matrix};1\right]\beta^{j}.$ (2.9)
Theorem 2.8.
Assume that $c,c-b\in\mathbb{C}\setminus\mathbb{Z}_{\leqslant 0}$ and $a,b,a-b\in\mathbb{C}\setminus\mathbb{Z}$. Set $\beta=-\frac{y}{x}$. Let $w>0$ be a number such that
$w>1+\max\left\{\Re(a-b),\Re(-b)+1\right\}$
and that the fractional parts of $w-\Re(a-b)$ and $w+\Re(b)-1$ are both in the interval $(\varepsilon,1)$, where $\varepsilon>0$ is small.
Then, under the conditions that $\frac{x-1}{x}y$ is bounded away from the points $b-a+k\,(k\in\mathbb{Z})$ and that
$x\to\infty,y\to\infty,\quad\left|\arg(1-x)\right|<\pi,\quad\left|\arg\left(\frac{1-x}{x}y\right)\right|<\pi,\quad 0<\beta_{1}\leqslant|\beta|\leqslant\beta_{2}<\infty,$
the function $\Phi_{1}\equiv\Phi_{1}[a,b;c;x,y]$ admits the asymptotic expansion
$\Phi_{1}={}$ $\frac{\Gamma(c)\Gamma(b-a)}{\Gamma(b)\Gamma(c-a)}A_{1}^{*}(x,y)+\frac{\Gamma(c)\Gamma(a-b)}{\Gamma(a)\Gamma(c-a)}\Bigl{(}A_{2}^{*}(x,y)+A_{3}^{*}(x,y)\Bigr{)}$
$+\frac{\Gamma(c)}{\Gamma(a)}A_{4}^{*}(x,y)+\mathcal{O}\left(\left|y\right|^{-\Re(b)-w}\right)+\mathcal{O}\left(\left|y\right|^{\Re(a-b-c)-N}\mathrm{e}^{\Re(y)}\right),$
where
$A_{1}^{*}(x,y)={}$ $\mathrm{e}^{-\beta}\left(1-x\right)^{-a}\sum_{k=0}^{M}\frac{\left(a\right)_{k}\left(c-b\right)_{k}}{\left(a-b+1\right)_{k}k!}\,_{2}F_{2}\left[\begin{matrix}1-b,c-a+k\\ c-b,a-b+1+k\end{matrix};\beta\right]\left(1-x\right)^{-k},$
$A_{2}^{*}(x,y)={}$ $\left(-\beta\right)^{b-a}\mathrm{e}^{-\beta}\left(1-x\right)^{-a}\sum_{k=1}^{M}\frac{\left(a\right)_{k}}{k!}\left(1-x\right)^{-k}\sum_{j=0}^{k-1}\frac{\left(c-a\right)_{j}\left(a-b\right)_{k-j}\left(a-c+1\right)_{k-j}}{j!\left(a\right)_{k-j}}\left(-\beta\right)^{j-k},$
$A_{3}^{*}(x,y)={}$ $\left(-\beta\right)^{b-a}\mathrm{e}^{-\beta}\left(1-x\right)^{-a}\sum_{k=0}^{M}\frac{\left(a\right)_{k}\left(c-a\right)_{k}}{\left.k!\right.^{2}}\,_{3}F_{3}\left[\begin{matrix}1,1-a,c-a+k\\ c-a,b-a+1,1+k\end{matrix};\beta\right]\left(1-x\right)^{-k},$
$A_{4}^{*}(x,y)={}$ $\beta^{a-c}\left(1-x\right)^{a-b-c}\mathrm{e}^{y}\sum_{k=0}^{N-1}a_{k}^{*}\left(x,y\right)y^{-k},$
with $M=\lfloor w+\Re(b-a)\rfloor\geqslant 1$, $N$ being any positive integer and $a_{k}^{*}(x,y)$ given by (2.9).
2.5.  Regime (vi): $x$ or $y$ small, $xy$ fixed
We first deduce the asymptotics of $\Phi_{1}\bigl{[}x,\frac{\eta}{x}\bigr{]}$ for large $x$.
Theorem 2.9.
Assume that $a,b\in\mathbb{C},\,c\in\mathbb{C}\setminus\mathbb{Z}_{\leqslant 0}$ and $a-b\in\mathbb{C}\setminus\mathbb{Z}$. Set $\eta=xy$. Then under the condition
$x\to\infty,\quad\left|\arg(-x)\right|<\pi,\quad 0<\eta_{1}\leqslant|\eta|\leqslant\eta_{2}<\infty,$
the function $\Phi_{1}\equiv\Phi_{1}[a,b;c;x,y]$ admits the asymptotic expansion
$\Phi_{1}=\left(-x\right)^{-a}\sum_{k=0}^{N-1}b_{k}^{(1)}(\eta)\,x^{-k}+\left(-x\right)^{-b}\sum_{k=0}^{N-1}b_{k}^{(2)}(\eta)\,x^{-k}+\mathcal{O}\left(\left|x\right|^{-\Re(a)-N}+\left|x\right|^{-\Re(b)-N}\right),$
where $N$ is any integer such that $N\geqslant\max\{1,|a|,|b|\}$, and
$b_{k}^{(1)}(\eta)$ $=\sum_{\begin{subarray}{c}m,n\geqslant 0\\ m+2n=k\end{subarray}}\frac{\left(a\right)_{m+n}\left(a-c+1\right)_{m}}{\left(a-b+1\right)_{m+n}m!\,n!}\eta^{n},$
$b_{k}^{(2)}(\eta)$ $=\sum_{\begin{subarray}{c}m,n\geqslant 0\\ m+n=k\end{subarray}}\frac{\left(b\right)_{m}\left(b-a+1\right)_{m-n}}{\left(b-c+1\right)_{m-n}m!\,n!}\eta^{n}.$
Proof..
The desired result follows immediately by first inserting $y=\eta x^{-1}$ into the series (2.3) and (2.4) and then expressing these double series as power series in $x^{-1}$. ∎
Now we establish the asymptotics of $\Phi_{1}\bigl{[}\frac{\eta}{y},y\bigr{]}$ for large $y$ by applying the uniformity approach, which was first employed in [13] and subsequently proposed in [12].
Theorem 2.10.
Assume that $a,b\in\mathbb{C}$ and $c\in\mathbb{C}\setminus\mathbb{Z}_{\leqslant 0}$. Set $\eta=xy$.
(1) Under the conditions that $y$ is bounded away from the points $a+\ell\,(\ell\in\mathbb{Z}_{\geqslant 0})$ and that
$y\to\infty,\quad\left|\arg(1-x)\right|<\pi,\,\left|\arg(-y)\right|<\pi,\quad 0<\eta_{1}\leqslant|\eta|\leqslant\eta_{2}<\infty,$
the function $\Phi_{1}\equiv\Phi_{1}[a,b;c;x,y]$ admits the asymptotic expansion
$\Phi_{1}={}$ $\frac{\Gamma(c)}{\Gamma(c-a)}\left(-y\right)^{-a}\sum_{k=0}^{N-1}c_{k}^{(1)}(\eta)\,y^{-k}+\frac{\Gamma(c)}{\Gamma(a)}y^{a-c}\mathrm{e}^{y}\sum_{k=0}^{N-1}c_{k}^{(2)}(\eta)\,y^{-k}$
$+\mathcal{O}\left(\left|y\right|^{-\Re(a)-N}+\left|y\right|^{\Re(a-c)-N}\mathrm{e}^{\Re(y)}\right),$
where $N$ is any integer such that $N\geqslant\max\{1,|a|,|a-c|\}$, and
$c_{k}^{(1)}(\eta)$ $=\left(-1\right)^{k}\sum_{\begin{subarray}{c}m,n\geqslant 0\\ m+2n=k\end{subarray}}\frac{\left(a\right)_{m+n}\left(a-c+1\right)_{m}\left(b\right)_{n}}{m!\,n!}\left(-\eta\right)^{n},$ (2.10)
$c_{k}^{(2)}(\eta)$ $=\sum_{\begin{subarray}{c}m,n\geqslant 0\\ m+n=k\end{subarray}}\frac{\left(c-a\right)_{m}\left(b\right)_{n}\left(1-a-n\right)_{m}}{m!\,n!}\eta^{n}.$ (2.11)
(2) Under the conditions that $y$ is bounded away from the points $a+\ell\,(\ell\in\mathbb{Z}_{\geqslant 0})$ and that
$y\to+\infty,\quad\left|\arg(1-x)\right|<\pi,\quad 0<\eta_{1}\leqslant|\eta|\leqslant\eta_{2}<\infty,$
the function $\Phi_{1}\equiv\Phi_{1}[a,b;c;x,y]$ admits the asymptotic expansion
$\Phi_{1}\sim\frac{\Gamma(c)}{\Gamma(a)}y^{a-c}\mathrm{e}^{y}\sum_{k=0}^{\infty}c_{k}^{(2)}(\eta)\,y^{-k}$
with the coefficients $c_{k}^{(2)}(\eta)$ given by (2.11).
Proof..
It suffices to establish Assertion (1), as the proof of (2) follows analogously through application of the estimate in [13, Theorem 2.6(ii)]. Now assume that the conditions stated in (1) are valid.
Recall the series representation [14, p. 75]
$\Phi_{1}[a,b;c;x,y]=\sum_{n=0}^{\infty}{}_{1}F_{1}\left[\begin{matrix}a+n\\ c+n\end{matrix};y\right]\frac{\left(a\right)_{n}\left(b\right)_{n}}{\left(c\right)_{n}}\frac{x^{n}}{n!},\quad|x|<1,|y|<\infty.$ (2.12)
Split the series into two parts, namely,
$\Phi_{1}\equiv\Phi_{1}[a,b;c;x,y]=\sum_{n=0}^{N-1}+\sum_{n=N}^{\infty}=:S_{1}+S_{2},$
where $N$ is any fixed positive integer. It remains to establish the asymptotic behaviors of $S_{1}$ and $S_{2}$.
Estimate on $S_{1}$. For $10\leqslant n\leqslant N-$ 1, recall the asymptotic expansion of ${}_{1}F_{1}$ [15, Eq. (5.8)]:
$\begin{split}{}_{1}F_{1}\left[\begin{matrix}a+n\\ c+n\end{matrix};y\right]\sim{}&\frac{\Gamma(c+n)}{\Gamma(c-a)}\sum_{m=0}^{\infty}\frac{\left(a+n\right)_{m}\left(1+a-c\right)_{m}}{m!}\left(-y\right)^{-a-n-m}\\ &+\frac{\Gamma(c+n)}{\Gamma(a+n)}\,\mathrm{e}^{y}\sum_{m=0}^{\infty}\frac{\left(1-a-n\right)_{m}\left(c-a\right)_{m}}{m!}y^{a-c-m},\quad y\to\infty.\end{split}$ (2.13)
Note that $\eta=xy$. Then substituting the first $N$ terms of both series in (2.13) into the definition of $S_{1}$ yields
$S_{1}={}$ $\frac{\Gamma(c)}{\Gamma(c-a)}\sum_{n=0}^{N-1}\frac{\left(a\right)_{n}\left(b\right)_{n}}{n!}\eta^{n}y^{-n}\sum_{m=0}^{N-1}\frac{\left(a+n\right)_{m}\left(1+a-c\right)_{m}}{m!}\left(-y\right)^{-a-n-m}$
$+\frac{\Gamma(c)}{\Gamma(a)}\sum_{n=0}^{N-1}\frac{\left(b\right)_{n}}{n!}\eta^{n}y^{-n}\mathrm{e}^{y}\sum_{m=0}^{N-1}\frac{\left(1-a-n\right)_{m}\left(c-a\right)_{m}}{m!}y^{a-c-m}$
$+\mathcal{O}\left(\left|y\right|^{-\Re(a)-N}+\left|y\right|^{\Re(a-c)-N}\mathrm{e}^{\Re(y)}\right)$
[…]
$+\mathcal{O}\left(\left|y\right|^{-\Re(a)-N}+\left|y\right|^{\Re(a-c)-N}\mathrm{e}^{\Re(y)}\right),$
where $c_{k}^{(1)}(\eta)$ and $c_{k}^{(2)}(\eta)$ are given by (2.10) and (2.11), respectively.
Estimate on $S_{2}$. For $n\geqslant N$, use the estimate [13, Theorem 2.5] to get
${}_{1}F_{1}\left[\begin{matrix}a+n\\ c+n\end{matrix};y\right]=\mathcal{O}\left(\left(n+1\right)^{\max\{0,\Re(c-a)\}}+\left(n+1\right)^{2|a-c|}\left|y\right|^{\Re(a-c)}\mathrm{e}^{\Re(y)}\right),$
which confirms that
$S_{2}=\mathcal{O}\left(\left|x\right|^{N}\left|y\right|^{\Re(a-c)}\mathrm{e}^{\Re(y)}\right)=\mathcal{O}\left(\left|y\right|^{\Re(a-c)-N}\mathrm{e}^{\Re(y)}\right).$
Combining the estimates on $S_{1}$ and $S_{2}$ gives the desired results. ∎
3.  Applications
In this section, we demonstrate the applicability of our results through two illustrative examples.
3.1.  Analytic continuations of $F_{M}$
Saran’s function $F_{M}$ is defined by ([19, Eq. (2.5)])
$\begin{split}F_{M}&\equiv F_{M}[\alpha_{1},\alpha_{2},\alpha_{2},\beta_{1},\beta_{2},\beta_{1};\gamma_{1},\gamma_{2},\gamma_{2};x,y,z]\\ &=\sum_{m,n,p=0}^{\infty}\frac{\left(\alpha_{1}\right)_{m}\left(\alpha_{2}\right)_{n+p}\left(\beta_{1}\right)_{m+p}\left(\beta_{2}\right)_{n}}{\left(\gamma_{1}\right)_{m}\left(\gamma_{2}\right)_{n+p}}\frac{x^{m}}{m!}\frac{y^{n}}{n!}\frac{z^{p}}{p!},\quad|x|+|z|<1,|y|<1,\end{split}$ (3.1)
where $\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}\in\mathbb{C}$ and $\gamma_{1},\gamma_{2}\in\mathbb{C}\setminus\mathbb{Z}_{\leqslant 0}$. Here we provide three analytic continuations for $F_{M}$.
Employing the series manipulation technique, we obtain an equivalent form of (3.1):
$\begin{split}F_{M}&[\alpha_{1},\alpha_{2},\alpha_{2},\beta_{1},\beta_{2},\beta_{1};\gamma_{1},\gamma_{2},\gamma_{2};x,y,z]\\ &\quad=\sum_{n=0}^{\infty}\frac{\left(\alpha_{2}\right)_{n}\left(\beta_{1}\right)_{n}}{\left(\gamma_{2}\right)_{n}}\,_{2}F_{1}\left[\begin{matrix}\alpha_{1},\beta_{1}+n\\ \gamma_{1}\end{matrix};x\right]{}_{2}F_{1}\left[\begin{matrix}\alpha_{2}+n,\beta_{2}\\ \gamma_{2}+n\end{matrix};y\right]\frac{z^{n}}{n!},\end{split}$ (3.2)
which, in view of (1.3) and the estimate [11, Eq. (2.2)]
$\left|{}_{2}F_{1}\left[\begin{matrix}a+n,b\\ c\end{matrix};z\right]\right|=\mathcal{O}\left(n^{-\min\{\Re(b),\Re(c-b)\}}\left(1+\left|1-z\right|^{-n}\right)\right),\quad n\in\mathbb{Z}_{\geqslant 1},n\to\infty,$
is absolutely convergent in the region
$\mathbb{D}_{M}:=\left\{(x,y,z)\in\mathbb{C}^{3}:|x|+|z|<1,|y|<\infty\right\}.$
Therefore, the series expansion (3.2) offers an analytic continuation of $F_{M}$ to the region $\mathbb{D}_{M}$. For more expansion formulae of $F_{M}$, we refer to [1, Section 3].
Substituting the integral
$\left(\beta_{1}\right)_{m+p}=\frac{1}{\Gamma(\beta_{1})}\int_{0}^{\infty}\mathrm{e}^{-t}t^{m+p+\beta_{1}-1}\mathrm{d}t,\quad\Re(\beta_{1})>0$
into (3.1) and swapping the order of integration and summation, we can get the Laplace integral for $F_{M}$:
$\begin{split}F_{M}&[\alpha_{1},\alpha_{2},\alpha_{2},\beta_{1},\beta_{2},\beta_{1};\gamma_{1},\gamma_{2},\gamma_{2};x,y,z]\\ &\quad=\frac{1}{\Gamma(\beta_{1})}\int_{0}^{\infty}\mathrm{e}^{-t}t^{\beta_{1}-1}\,_{1}F_{1}\left[\begin{matrix}\alpha_{1}\\ \beta_{1}\end{matrix};xt\right]\Phi_{1}\left[\alpha_{2},\beta_{2};\gamma_{2};y,zt\right]\mathrm{d}t,\end{split}$ (3.3)
which corrects Saran’s original form [20, p. 134, Eq. (4)]. Our results in Section 2.3 regarding asymptotics of $\Phi_{1}[y,zt]$ for large $t$ then provide a sufficient convergence condition for (3.3):
$\Re(x+z)<1,\ \Re(\beta_{1})>0.$
Hence the integral (3.3) gives the second continuation of $F_{M}$.
The final continuation of $F_{M}$ follows from the following Mellin-Barnes integral, which is established by repeating the analysis in the proof of Theorem 2.3 and applying the asymptotic results of the Gauss hypergeometric function ${}_{2}F_{1}$ for large parameters in [11, Section 2.1].
Theorem 3.1.
Let $\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}\in\mathbb{C},\,\gamma_{1},\gamma_{2}\in\mathbb{C}\setminus\mathbb{Z}_{\leqslant 0}$ and
$\mathbb{V}_{F_{M}}=\left\{(x,y,z)\in\mathbb{C}^{3}:\begin{array}[]{c}x\neq 1,\ y\neq 1,\ \left|\arg(1-x)\right|<\pi,\ \left|\arg(1-y)\right|<\pi,\\ z\neq 0,\ \left|\arg(-z)\right|<\pi,\\ \left|\arg(1-x)+\arg(-z)\right|<\pi\end{array}\right\}.$
Then for $(x,y,z)\in\mathbb{V}_{F_{M}}$,
$F_{M}$ $[\alpha_{1},\alpha_{2},\alpha_{2},\beta_{1},\beta_{2},\beta_{1};\gamma_{1},\gamma_{2},\gamma_{2};x,y,z]$
$\quad=\frac{1}{2\pi\mathrm{i}}\frac{\Gamma(\gamma_{2})}{\Gamma(\alpha_{2})\Gamma(\beta_{1})}\int_{\mathfrak{L}_{\mathrm{i}\sigma\infty}}{}_{2}F_{1}\left[\begin{matrix}\alpha_{1},\beta_{1}+s\\ \gamma_{1}\end{matrix};x\right]{}_{2}F_{1}\left[\begin{matrix}\alpha_{2}+s,\beta_{2}\\ \gamma_{2}+s\end{matrix};y\right]\frac{\Gamma(\alpha_{2}+s)\Gamma(\beta_{1}+s)}{\Gamma(\gamma_{2}+s)}\Gamma(-s)\left(-z\right)^{s}\mathrm{d}s,$
where the path $\mathfrak{L}_{\mathrm{i}\sigma\infty}$, starting at $\sigma-\mathrm{i}\infty$ and ending at $\sigma+\mathrm{i}\infty$, is a vertical line intended if necessary to separate the poles of $\Gamma(\alpha_{2}+s)\Gamma(\beta_{1}+s)$ from the poles of $\Gamma(-s)$.
3.2.  $1D$ Glauber-Ising model
In 2000, Godrèche and Luck [10] proved that, the two-time auto-correlator in the $1D$ Glauber-Ising model quenched to some temperature $T>0$, starting from a fully disordered initial configuration (or infinite temperature), is expressed in terms of the Humbert function $\Phi_{1}$ (see [10, Eq. (4.22)]):
$C_{0}(s+\tau,s)=\frac{2}{\pi}\sqrt{\frac{2s}{\tau}}\,\mathrm{e}^{-\frac{1}{2}\mu^{2}\tau}\,\Phi_{1}\left[\frac{1}{2},1;\frac{3}{2};-\frac{2s}{\tau},-\mu^{2}s\right].$
Here the two times are $s$ and $t=s+\tau$, the scaling variables are $2s/\tau$ and $\mu^{2}s=2s/\tau_{\text{eq}}$, and the parameter $\mu$ is related to temperature via the identity [10, Eq. (1.8)]
$\mu\equiv-\log\,\tanh(\beta J)=-\log\,\tanh\left(k_{\text{Boltz}}TJ\right),$
where $k_{\text{Boltz}}$ is the Boltzman constant.
We proceed to give a rigorous mathematical proof of the results stated in Assertions (i) and (ii) of [10, Section 4]. Now let us present an alternative formulation of Assertions (i) and (ii):
  • Assertion (i): At zero temperature, the autocorrelation function simplifies to
    .$\bigl{.}C_{0}(s+\tau,s)\bigr{|}_{\mu=0}=\frac{2}{\pi}\arctan\sqrt{\frac{2s}{\tau}}.$ (3.4)
  • Assertion (ii): For $s\to\infty$ with $\tau$ and $\mu$ fixed, the autocorrelation function converges to the equilibrium
    $C_{0,\text{eq}}(\tau)=\text{erfc}\left(\sqrt{\frac{\tau}{\tau_{\text{eq}}}}\right),$ (3.5)
    where erfc denotes the complementary error function.
Proof of Assertions.
Write $x=2s/\tau$ and $y=\mu^{2}s=2s/\tau_{\text{eq}}$, and denote
$F(x,y)\equiv C_{0}(s+\tau,s)=\frac{2}{\pi}\sqrt{x}\,\mathrm{e}^{-\frac{y}{x}}\Phi_{1}\left[\frac{1}{2},1;\frac{3}{2};-x,-y\right].$
(i) When $\mu=0$, we have $y=0$ and thus obtain from (1.1) that
$F(x,0)=\frac{2}{\pi}\sqrt{x}\,\Phi_{1}\left[\frac{1}{2},1;\frac{3}{2};-x,0\right]=\frac{2}{\pi}\sqrt{x}\,\,_{2}F_{1}\left[\begin{matrix}1/2,1\\ 3/2\end{matrix};-x\right],$
which, by using the reduction formula [18, Eq. (15.4.3)]
${}_{2}F_{1}\left[\begin{matrix}1/2,1\\ 3/2\end{matrix};-z^{2}\right]=z^{-1}\arctan z,$
yields the equivalent statement of (3.4):
$F(x,0)=\frac{2}{\pi}\arctan\sqrt{x}=\frac{2}{\pi}\arctan\sqrt{\frac{2s}{\tau}}.$
(ii) As $s\to\infty$, we have $x\to+\infty$ and $y\to+\infty$. To apply Theorem 2.8, we have to impose an extra restriction:
$\frac{x+1}{x}y\text{ is bounded away from the half-integers }k-\frac{1}{2}\,(k\in\mathbb{Z}_{\geqslant 1}).$ (3.6)
Then as $s\to\infty$ with (3.6) valid, Theorem 2.8 implies for $\rho=\frac{y}{x}$ that
$\Phi_{1}\left[\frac{1}{2},1;\frac{3}{2};-x,-y\right]\sim\Gamma\left(\frac{3}{2}\right)\Gamma\left(\frac{1}{2}\right)\mathrm{e}^{\rho}x^{-\frac{1}{2}}+\frac{\Gamma\left(\frac{3}{2}\right)\Gamma\left(-\frac{1}{2}\right)}{\Gamma\left(\frac{1}{2}\right)}\rho^{\frac{1}{2}}\mathrm{e}^{\rho}x^{-\frac{1}{2}}\,_{1}F_{1}\left[\begin{matrix}1/2\\ 3/2\end{matrix};-\rho\right].$
Employing the identity [18, Eq. (13.6.7)]
${}_{1}F_{1}\left[\begin{matrix}1/2\\ 3/2\end{matrix};-z^{2}\right]=\frac{\sqrt{\pi}}{2z}\bigl{(}1-\text{erfc}(z)\bigr{)},$
we have
$\Phi_{1}\left[\frac{1}{2},1;\frac{3}{2};-x,-y\right]$ $\sim\Gamma\left(\frac{3}{2}\right)\Gamma\left(\frac{1}{2}\right)\mathrm{e}^{\rho}x^{-\frac{1}{2}}+\frac{\sqrt{\pi}\,\Gamma\left(\frac{3}{2}\right)\Gamma\left(-\frac{1}{2}\right)}{2\,\Gamma\left(\frac{1}{2}\right)}\mathrm{e}^{\rho}x^{-\frac{1}{2}}\bigl{(}1-\text{erfc}(\sqrt{\rho})\bigr{)}$
$=\frac{\pi}{2}x^{-\frac{1}{2}}\mathrm{e}^{\rho}\cdot\text{erfc}(\sqrt{\rho}).$
Accordingly, we get the desired limit
$\lim_{s\to\infty}F(x,y)=\text{erfc}(\sqrt{\rho})=\text{erfc}\left(\sqrt{\frac{\tau}{\tau_{\text{eq}}}}\right).\qed$
4.  Discussion
We derived full asymptotic expansions of the Humbert function $\Phi_{1}$ in various limiting cases of its variables. The conditions in Theorems 2.7-2.10 can be relaxed through careful adjustments of the remainder terms, extending their validity to the parameter ranges $0<|\beta|\leqslant\beta_{0}<\infty$ and $0<|\eta|\leqslant\eta_{0}<\infty$. However, the asymptotic expansions in Theorems 2.8 and 2.10 are restrictive in the sense that the variables must remain bounded away from a countable set of exceptional values. In view of the series representations (2.2) and (2.12), the restrictions on exceptional values may be removed by deducing the asymptotic expansions of
${}_{1}F_{1}\left[\begin{matrix}a\pm n\\ b\pm n\end{matrix};z\right],\quad z\to\infty,$
which hold uniformly with respect to $n\in\mathbb{Z}_{\geqslant 0}$. More generally, it is necessary to extend the asymptotic expansions for ${}_{1}F_{1}[a;b;z]$ established in [24, Theorem 3.1], which reads
$\frac{\Gamma(a)}{\Gamma(b)}z^{b-a}\mathrm{e}^{-z}\,_{1}F_{1}\left[\begin{matrix}a\\ b\end{matrix};z\right]\sim\mathrm{e}^{-z\mathcal{A}}f_{0}(\mu)\sum_{n=0}^{\infty}\widetilde{f_{n}}(a)z^{-n},\quad z\to\infty$
and requires at least one of the parameters $a,b$ and $z$ is large, to arbitrary complex values of $a,b$ and $z$.
We also established three analytic continuations of Saran’s function $F_{M}$. Notably, $F_{M}$ appears to have finer properties compared to Saran’s $F_{K}$ function. The fundamental properties of $F_{M}$ can be systematically derived by adapting techniques from prior studies on $F_{K}$ [6, 16, 17].
Finally, we would like to emphasize that all methods employed in this paper (see Table 1) are based on the series representations of multivariate special functions, rather than the integral representations. Among these methods, the uniformity approach currently has significant limitations (see [12, Section 7]). Now we propose three potential ways to overcome them:
  1. 1.
    Derive asymptotic expansions for the generalized hypergeometric function ${}_{p}F_{q}$ when at least one of the parameters and the variable $z$ becomes large, as illustrated by the ${}_{1}F_{1}$ case discussed above.
  2. 2.
    Deduce asymptotic expansions for the ratios of the gamma functions
    $\frac{\Gamma(z+\lambda_{1}(n,z))\Gamma(z+\lambda_{2}(n,z))\cdots\Gamma(z+\lambda_{p}(n,z))}{\Gamma(z+\mu_{1}(n,z))\Gamma(z+\mu_{2}(n,z))\cdots\Gamma(z+\mu_{q}(n,z))},\quad z\to\infty$
    valid uniformly for $n\in\mathbb{Z}_{\geqslant 0}$, where the $\lambda_{j}(n,z)$’s and $\mu_{k}(n,z)$’s are quantities involving $n$ and $z$.
  3. 3.
    Explore high-dimensional generalizations of the uniformity approach.
These ways each merit further in-depth investigation.
Acknowledgement
We wish to thank Malte Henkel for useful discussions with the first author.